OpenStudy (anonymous):

Robert is entering a contest at a local radio station. all of the cards with the correct answer will be placed in a barrel, and from it will be drawn the name of the winner. Robert submits 200 post cards with the correct answer for the drawing. If 250,000 entries were received and 16,000 of them are correct, what is the probability that Robert will win?

OpenStudy (anonymous):

Student government sold 500 tickets for a raffle. They sell each ticket for $1. An officer draws one winning ticket from a box. If you've purchased 10 tickets, what is the probability that you'll win?

OpenStudy (anonymous):

@younglove

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

@sammie2017

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

Robert's chances 200/16000

OpenStudy (triciaal):

200/16000 when simplified = 1/80

OpenStudy (triciaal):

probability of same name = 1/10 * 10/20

OpenStudy (triciaal):

correction not same name but own name own name = 1/20

OpenStudy (triciaal):

Raffle 10 tickets out of 500 tickets = 1/50 chance

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

The gardener mixed up the petunia he bought at the plant shop. None of them are blooming, but he knows he bought 14 purple, 8 red, and 4 white. If he just selects one at random, what is the probability it is red

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

@bibby @triciaal

OpenStudy (anonymous):

@zBrandz23

OpenStudy (studygurl14):

14 purple 8 red 4 white 14 + 8 + 4 =26 total number of red/total = prob of picking red 8 / 26 simiplify 4/13

OpenStudy (anonymous):

Thank You!

OpenStudy (studygurl14):

yw

OpenStudy (anonymous):

The bank is giving away free car blankets to every one who receives a car loan through the bank. The bank now has only a couple of blankets left in each color. There are 2 red, 1 blue, and 8 yellow. Sue has just received an approved car loan. What is the probability the bank will give her a blue blanket.

OpenStudy (anonymous):

wouldn't it be 1/11

OpenStudy (anonymous):

1 blue/11 over all colors

OpenStudy (anonymous):

@dragonmaster917

OpenStudy (studygurl14):

2 red 1 blue 8 yellow blue blanket/total blankets So yes, 1/11

OpenStudy (anonymous):

Thanks for the re-assured answer =)

OpenStudy (anonymous):

I also had two other questions i wasn't completely confident about could you help me with those last 2

OpenStudy (studygurl14):

Sure, could you post them *here* http://openstudy.com/study#/updates/54381a69e4b0c20f78101d9f

OpenStudy (studygurl14):

i'm trying to build my popularity and fan base^

OpenStudy (bibby):

I was one of the first maaaan

OpenStudy (anonymous):

Bill would like to use his computer at work to do a budget statement. He knows that 10 of the computers do not work, 3 of the computers have virus that destroys your documents, and 9 of the computers are working. What is the probability that the 1st computer bill tries works correctly. I guessed 9/22

OpenStudy (anonymous):

lol u guys but yea 9/22 ? Does it seem accurate

OpenStudy (anonymous):

does that medal some one gave me mean yes lol

OpenStudy (studygurl14):

10 do not work 3 virus that destroys documents (basically means it doesn;t work) 9 work correctly so... 9/22 is correct!

OpenStudy (bibby):

9/10+3+9=9/22

OpenStudy (bibby):

yeah

OpenStudy (anonymous):

i didn't know you could get medals for freely runnin your mouth j/p j/p

OpenStudy (bibby):

that's the only reason I have 92 smartscore brah

OpenStudy (anonymous):

lmaoooo @bibby and thx @StudyGurl14 you a genius and i'm steadily workin on being great in math too, anyhow just got one more question i'd like someone to check if if they dnt mind

OpenStudy (bibby):

just post it instead of trying to not step on our toes :p

OpenStudy (studygurl14):

ok

OpenStudy (anonymous):

Amelia sets her cd player to randomly select songs from the 4 cds she's loaded. Disc a has 14 songs, Disk b 9, c 11, and d 10, what is the probability that first song she hears is disk d

OpenStudy (anonymous):

10/34 5/17 is what I got

OpenStudy (studygurl14):

14 disk a 9 disk b 11 disk c 10 disk d so...*calculating* 10/44 Is the answer, non-simplified